Find the center of the conic 9x²-4y²-18x-24y-63=0
Answer by lwsshak3(11628)

Find the center of the conic 9x²-4y²-18x-24y-63=0
**
9x²-4y²-18x-24y-63=0
complete the square
9(x^2-2x+1)-4(y^2+6y+9)=63+9-36=36
9(x-1)^2-4(y+3)^2=36
(x-1)^2/4-(y+3)^2/9=1
This is an equation of a hyperbola with horizontal transverse axis of the standard form:
(x-h)^2/a^2-(y-k)^2/b^2=1, with (h,k) being the (x,y) coordinates of the center.
For given equation:
center: (1,-3)